PLEASEEE
1. Find functions f and g so that f [g(x)] = H(x) where H(x)=(1+x^2)^3
2. If g(x)=x^2+4x+2 find g(2) and g(x+2)
3. For f(x)=4x+1 and g(x)=x^2-2x-4 find (g-f)(x). Simplify the answer.
4.What is the function which would intersect the y-axis 2 positions up(vertical) from the function f(x)=x^2
5. Using the graphing utility to approximate any relative minimum or maximum points of f(x)=x^2-6x.
THANK YOU VERY MUCH

Answer by stanbon(48535)

1. Find functions f and g so that f [g(x)] = H(x) where H(x)=(1+x^2)^3
g(x) = 1+x^2
f(x) = x^3
---------------------------
2. If g(x)=x^2+4x+2 find g(2) and g(x+2)
g(2) = 1+2^2 = 5
----
g(x+2) = 1 + (x+2)^2 = x^2 + 4x +5
==========================
3. For f(x)=4x+1 and g(x)=x^2-2x-4 find (g-f)(x). Simplify the answer.
(g-f)(x) = x^2-2x-4 -(4x+1) = x^2 -6x -5
-------------------

4.What is the function which would intersect the y-axis 2 positions up(vertical) from the function f(x)=x^2
f(x) = x^2 +2
---------------------
5. Using the graphing utility to approximate any relative minimum or maximum points of f(x)=x^2-6x.
(3,-9)